40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 132 of 354  Not logged in ELOG logo
ID Date Author Type Category Subject
  11217   Tue Apr 14 11:19:52 2015 SteveUpdatePEMlarge chamber has arrived

It is here.

Quote:

The 40m fenced area will start storing this large ~ 8000 lbs chamber on April 14. The asphalt will be cut, jack hammered the next 2-3 days in order to lay concrete.

Their schedule is from 8 to 5 starting tomorrow.  We are asking them to work from 6 to 3pm

ETMX is about 12-15 ft away

 

Attachment 1: ETMXfriend.png
ETMXfriend.png
Attachment 2: 8000lbs.png
8000lbs.png
  11216   Mon Apr 13 19:34:02 2015 ericqUpdateIOOModulation Frequency Tuned to IMC Length

I've been fiddling with the mode cleaner and green beat box today, to try and get an absolute frequency calibration for MC2 motion. The AC measurements have all turned out weird, I get fractional power laws instead of the 1/f^2 that we expect from the MC2 pendulum. At DC, I get a rough number of 15 green kHz per MC2 count, but this translates to ~7e-10 m/count which is in contrast to the 6e-9 m/count from 2009. I will meditate on this a bit. 


In any case, while working at the IOO rack, I tuned the 11MHz modulation frequency, as was done in ELOGs 9324 and 10314, by minimizing one of the beats of the 11MHz and 29.5MHz sidebands. 

The new modulation frequency / current IMC FSR is 11.066209 +- 1 Hz, which is a only a few ppm change from the tuning from last July.

This implies a IMC round trip length of 27.090800m +- 2um.

Attached is a plot showing the beat of 55-29.5 going down as I changed the marconi frequency. 

 

Attachment 1: fMod_tuning.pdf
fMod_tuning.pdf
  11215   Fri Apr 10 18:39:39 2015 ericqUpdateLSCRelative ETM calibration (Rough MC2 calibration)

I didn't verify that the loop gain was low enough at the excitation frequencies. blush

I put a 1kHz ELP in both arm servos, and got cleaner data for both. The ETM numbers are pretty much consistent with the previously posted ones, and the MC2 data now is consistent across frequencies. However, the MC2 numbers derived from each arm are not consistent.

Now:

  • ETMX / ITMX: 2.831 +- 0.043
  • MC2 / ITMX: 3.260 +- 0.062
  • ETMY / ITMY: 2.916 +- 0.041
  • MC2 / ITMY: 3.014 +- 0.036

With the data from ELOG 8242, this implies:

  • ETMX: 13.31 +- 0.21 x 10-9 / f2 m/counts
  • ETMY: 13.59 +- 0.20 x 10-9 / f2 m/counts
  • MC2 in Xarm meters : 15.32 +- 0.30 x 10-9 / f2 m/counts
  • MC2 in Yarm meters : 14.04 +- 0.18 x 10-9 / f2 m/counts
This is, of course, pretty fishy. Each arm sees the same frequency fluctuation of the light coming out of the IMC, especially given that the MC2 to arm data was taken simultaneously for both arms. Now, one possible source of this kind of mismatch would be a mismatch of the arm lengths, but there is no way they differ by 10%, as they would have to in order to explain the above numbers. To me, it seems more likely that the ITM calibrations are off. 
Attachment 1: betterCal.png
betterCal.png
  11214   Fri Apr 10 17:05:45 2015 ericqUpdateLSCRelative ETM calibration (Rough MC2 calibration)

I did a quick measurement get an idea of the ETM actuator calibration, relative to the ITMs. This will still hold if/when we revisit the ITM calibration via the Michelson. 

For the test masses, I locked the arms individually using MC2 as the actuator, and took transfer functions from the SUS-[OPTIC]_LSC_EXC point to the PO[X/Y]_I_ERR error signals. There were two points with coherence less than 99% that I threw away. I then took the fraction at each point, and am using the standard deviation of those fractions as the reported random error, since the coherence was super high for all points, making the error of each point negligible relative to their spread. 

This gives:

  • ETMX/ITMX: 2.765 +- 0.046
  • ETMY/ITMY: 2.857 +- 0.029

With the data from ELOG 8242, this implies:

  • ETMX: 13.00 +- 0.22 x 10 -9 / f2 m/counts
  • ETMY: 13.31 +- 0.15x 10 -9 / f2 m/counts

MC2 data was taken with the arms locked with the ETMs. The results are not so clean, the fractions don't line up; there is some trend with excitation frequency... The ratio is around the same as the ETMs, but I'm not going to quote any sort of precision, since I don't fully understand what's happening. Kind of a bummer, because it struck me that we could get an idea of the arm length mismatch by the difference in IMC frequency / arm FSR. I'll think about this some more...

Attachment 1: quickCal.png
quickCal.png
  11213   Fri Apr 10 12:09:19 2015 ericqUpdateLSCSome small progress, may have DAC problem?
Quote:

At the very end, the last 10 seconds or so, the POP110 power goes down, and sits at about half it's maximum value.  POP22 isn't quite as bad, in that it still touches the max, but the RIN is about 50%.  The carrier DC signals (TRX, TRY, POPDC) don't see this huge jump.  I don't think I was touching anything the last few tens of seconds.  I'm not sure yet how I can so significantly lose sideband power, without losing a similar amount of carrier power. 

I saw this same kind of behavior in my locklosses on Wednesday night; we should check out the 165 data, and see if the 3f PRCL error signal shows some drift away from zero.

Also, it's odd that CARM_IN1 and REFL11_I_ERR have different low frequency behavior in the plot you posted. I guess they have some difference in demodulation phase.  REFL11_I's bump at -40sec coincides with the dip in arm power and a rise in REFLDC, but ASDC seems pretty smooth, so maybe it is a real CARM fluctuation.

I set the REFL11 analog demodulation angle (via cable length) about a year ago (ELOG 9850), with some assumption about PRCL having the same demod angle as CARM, but this was probably set with the arms misaligned. We should recheck this; maybe we're coupling some other junk into CARM. 

  11212   Fri Apr 10 03:44:51 2015 JenneUpdateLSCSome small progress, may have DAC problem?

Small steps tonight, but all in the forward direction.

On one of my better locks, I saw a kind of weird phenomenon with the PRMI sideband powers versus the carrier powers:

For the last 100 seconds of this plot, I'm all 1f.  Alignment is being handled mostly by Q's DC coupled ITM oplevs, and the transmission QPD ASC loops, although I was trying to adjust the offsets in the ASC loops to improve transmission for a bit.

At the very end, the last 10 seconds or so, the POP110 power goes down, and sits at about half it's maximum value.  POP22 isn't quite as bad, in that it still touches the max, but the RIN is about 50%.  The carrier DC signals (TRX, TRY, POPDC) don't see this huge jump.  I don't think I was touching anything the last few tens of seconds.  I'm not sure yet how I can so significantly lose sideband power, without losing a similar amount of carrier power. 

The ring-ups at about -70sec in the CARM and DARM outs are the bounce mode. 

I tried looking at 2D histograms of different combinations of channels, for the time around -30 seconds where things looked pretty clean.  It looks like the offsets that Q put in last night (+1 for MICH_B and -3 for PRCL_B) are still about right.  The PRCL_IN1 and MICH_IN1 were centered around zero at the maximum power points.  CARM and DARM had small offsets, which I put into the DARM_B and CARM_B filter banks (0.0066 for DARM_B, and 0.027 for CARM_B), although these are small enough that I don't know that they really do anything.


As a break from locking for a little while, I tried to see if I could get the TT3 and TT4 DAC channels to work for me.  I had hoped it would be a quick and easy task, but I'm not seeing signal out.  Since it wasn't working, I decided to go back to locking for the night, and look into the DAC in the daytime.  I want to use one channel as the IN2 input of the CM board, and another as the external modulation input to the Marconi for transfer functions, so I need them to work. 

As a side note on the input to the Marconi situation, it occurred to me that instead of laying a new cable, I can borrow the POP55 heliax.  We don't have a POP55 diode right now, and the other end comes out across the hall from the Marconi, so it would be pretty easy to have a medium-length cable go from ITMX table to the Marconi.  Objections to this? 

Attachment 1: PRMI_sb_powers_9Apr2015.png
PRMI_sb_powers_9Apr2015.png
  11211   Thu Apr 9 07:41:22 2015 SteveUpdatePEMjackhammering at ETMX

There was a period of short unexpected jackhammering this morning. I asked them to stop.

The good mood of GTRX was not changed.smiley

 

Attachment 1: moreJackHam.png
moreJackHam.png
Attachment 2: GTRX.jpg
GTRX.jpg
  11210   Thu Apr 9 02:58:26 2015 ericqUpdateLSCAll 1F, all whitened

blarg. Chrome ate my elog. 

112607010 is the start of five minutes on all whitened 1F PDs. REFL55 has more low frequency noise than REFL165, I think we may need more CARM supression (i.e. we need to think about the required gain). This is also supported by the difference in shape of these two histograms, taken at the same time in 3f full lock. The CARM fluctuations seem to spread REFL55 out much more.  

I made some filters and scripts to do DC coupling of the ITM oplevs. This makes maintaining stable alignment in full lock much easier. 

I had a few 15+ minute locks on 3f, that only broke because I did something to break it.  

Here's one of the few "quick" locklosses I had. I think it really is CARM/AO action, since the IMC sees it right away, but I don't see anything ringing up; just a spontaneous freakout. 

Attachment 1: quickLockLoss.png
quickLockLoss.png
Attachment 2: 55_1.png
55_1.png
Attachment 3: 55_2.png
55_2.png
  11209   Wed Apr 8 21:10:55 2015 manasaUpdateGeneralWaking up the PDFR measurement system

I was poking around with the PDFR hardware today.

I moved the Agilent which had its screen projected on the monitor. I have put it back...but please verify the settings before using it for tonight.

  11208   Wed Apr 8 13:26:47 2015 ericqUpdateLSCREFL55 signal back to its normal ADC inputs

As the POP55 demod board is actually demodulating the REFL55 signal, I have connected its outputs to the REFL55 ADC inputs. Now, we can go back to using the REFL55 input matrix elements, and the data will be recorded. 

I have changed the relevant lines in the locking script to reflect this change. 

  11207   Wed Apr 8 03:44:48 2015 JenneUpdateLSCMediocre locking night

It was only a mediocre locking night.  I was foiled for a long time by 2 things, both of which are now taken care of in the scripts, so I don't waste so much time again.

  • Somehow FM4 (LP700) in the CARM_A filter bank was turned off.  It took me a long time to figure this one out. 
    • At first, I had a 2056Hz oscillation, and then if I notched that, I would have problems at the edges of my notch.
    • I turned on the LP1000 in CARM_B (which we don't usually use) to fight the 2k resonances, but then I had violin mode problems. 
    • I couldn't turn off the violin mode filters on MC2 for the transition, so the edges of these notches were causing instability.
    • Anyhow, in the end, I realized that FM4 of CARM_B wasn't on, but it usually is.
    • It is now turned on in the carm_cm_up.sh script.
  • After that fiasco, I had trouble turning on the ASC loops.
    • Turns out we had left the offsets in place from last night in the ASC loops, so that when I zeroed the outputs of the transmission QPDs, the offsets in the ASC loops didn't make any sense, and they pushed the IFO severely out of alignment.
    • Now the ASC down script (which is run by the carm down script) zeros the filter bank offset values

Scripts are checked into the svn.


I used Q's handy-dandy 2D histogram plotter (..../scripts/general/dataHist, which I have taken the liberty of adding to the svn) to set the PRCL offset when I was locked on REFL165.  Here is a version of the plot, when I had an offset of +10 in the PRCL filter bank. There was so much noise on the PRCL input that I quit bothering to try and put in an excitation or ramp the offset value.  Note that I have since moved this offset to PRCL_A's offset instead, so taking this plot again should have PRCL_IN1 centered around zero.

I had trouble doing something similar for PRCL when I was locked on REFL55.  At first, the offset was so poor that POP110 was only about half the value it was when locked on REFL165, and it had a huge amount of RIN.  I tried just doing a z avg of the PRCL_B_IN1 (REFL55I) while locked on REFL165, and that said that REFL55I had an offset of +33.8 counts, so I tried an offset of -33.8 counts to get to zero.  But, that was still terrible for POP110 power.  As I increased the offset, eventually up to +30 counts, POP110 kept getting better and better.  I lost lock at that point (while trying to get 10 sec of histogram data), so I'm not sure that +30 is the final value.  I want to also get equivalent histograms with POP22 and POPDC (and maybe arm transmissions?) as the X-axis on these plots.  There's no excitation, so all of these can be collected at once.


By babysitting the ITM alignment (looking at the rough DC values of the optical lever error siganls), and doing a little adjustment of the ASC differential offset, I was able to keep lock a few times for more than 2 or 3 minutes while all 1f.  Not a whole lot longer than that though, even if I wasn't "poking" the interferometer other than maintaining alignment.

Attachment 1: PRCL_R165_offsetplus10.png
PRCL_R165_offsetplus10.png
  11206   Tue Apr 7 04:21:45 2015 ericqUpdateASCAngular Control during Locking

[J, Q]

Alignment is making it tough for locks to last more than 10 minutes. Many (but not all) locklosses correlate with some optic drifting away, and taking all of the light with it. The other locklosses are the quick ones that seem to pop up out of nowhere; we haven't made any headway on these. We wanted to get to a state where we could just let the interferometer sit for some minutes, to explore the data, but got caught up with alignment and PRMI things.

We're finding that both ITMs experience some DC force when entering full PRFPMI lock. I will calculate the torque expected from radiation pressure + offset beam spot, especially for ITMX, where we choose the spot position to be uncontrolled by ASS. 

I set up the QPD ASC servos to act in a common/differential way on the ETMs. The C1:ASC-XARM_[PIT/YAW] filter modules act on the common alignment, whereas the C1:ASC-YARM_[PIT/YAW] filter modules act on the differential alignment. This can soon be cleaned up with some model renaming to reduce confusion. 

Using DC oplev values as a guide, we are hand tuning ITM alignment once the AO path is engaged and we see the DC drift occurring. Then, we set the QPD servo offsets and engage them. 

In this manner, we were able to lock the interferometer at:

  • Arm transmission 150 x single arm power
  • POPDC indicated a recycling gain of ~5.5
  • ASDC/POPDC indicated a contrast of 99.8%
  • REFLDC indicated a visibility of 80%

We made the PRMI transition to 1f numerous times, but found that the sideband power fluctuations would get significantly worse after the transition. 

We found that the gains that were previously used were too small by a factor of a few. There is a DC change visible in REFL165 before and after the transition (Also POP55, aka REFL55, is not DQ'd angry). Really, it isn't certain that we've zero'd the offset in the CARM board either, so REFL55's zero crossing isn't necessarily more trustworthy that REFL165's. We can go back in the data and do some 2D histograming to see where in the error signal space the sideband power is maximized. 

Jenne reports:

  • The all RF transition succeeded 13/29 times. 
  • PRMI 1f transision succeeded 10/10 times. 
  11205   Tue Apr 7 04:17:42 2015 JenneUpdatePEMjackhammering at ETMX

Q is writing the locking elog for the night, but just to reply to this thread:  The IFO worked well tonight, so things are at least not broken.

Quote:

Unfortunately, this kind of trend plot is not detailed enough to know if something has gone bad in a quantitative way. But at least we can tell that the suspension wire didn't break.

 

  11204   Mon Apr 6 13:23:40 2015 ranaUpdatePEMjackhammering at ETMX

Unfortunately, this kind of trend plot is not detailed enough to know if something has gone bad in a quantitative way. But at least we can tell that the suspension wire didn't break.

  11203   Mon Apr 6 09:13:49 2015 SteveUpdatePEMjackhammering at ETMX
Quote:

The 40m fenced area will start storing this large chamber on April 14. The asphalt will be cut, jack hammered the next 2-3 days in order to lay concrete.

Their schedule is from 8 to 5 starting tomorrow.  We are asking them to work from 6 to 3pm

ETMX is about 12-15 ft away

Jackhammering was happening around 7:30am

It looks like it did no harm. It is too early to say what may have moved. Rana's worrisome  email was late.

The ground preparation is completed.

Attachment 1: jackhammerCalb.png
jackhammerCalb.png
Attachment 2: jackHcalETMX.jpg
jackHcalETMX.jpg
Attachment 3: jackHetmx.png
jackHetmx.png
  11202   Sat Apr 4 17:54:03 2015 ranaUpdateModern ControlPreliminary PRMI angular Wiener results

This is a very cool result. I'm surprised that it can work so good. Please post what frequency dependent weighting you used on the target before running the Wiener code.

I think its important to tune it to keep the low frequencies from getting amplified by a factor of 10 (as they are in your plots). The seismometers are all just noise acting like thermometers and tilt sensors below 0.1 Hz. Temperature and tilt do not couple to our interferometer very much.

It also seems weird that you would need 20th order filters to make it work good. In any case, you can always split the SOS up into pieces before making the digital filters. For LLO, we used 3 buttons in some cases.

Quote:

Before locking for the evening, I wanted to try again implementing the Wiener filters that I had designed back in Jaunary (elog 10959). 

  11201   Fri Apr 3 19:35:14 2015 JenneUpdateSUSBS oplev centered

I think that this happens when the beam gets too close to the edge of the QPD.  We see this regularly in the ETMs, if they've been kicked a bit, but not enough to trip the watchdogs.  I think it might be the step/impulse response of the RES3.3 filter, which rings for almost 20 seconds. 

Anyhow, I've just recentered the BS oplev.  It was at -21urad in pitch, and had more than 400 counts on the top two quadrants, but only about 100 counts on the bottom two.  Now it's around 300 counts on all 4 quadrants.

As a totally unrelated aside, I have installed texlive on Donatella, so that I could run pdflatex.

  11200   Fri Apr 3 15:15:55 2015 SteveUpdateSUSBS oplev

I saw this kicking before  

Quote:

The BS oplev has been misbehaving and kicking the optic from time to time since noon. The kicks are not strong enough to trip the watchdogs (current watchdog max counts for the sensors is 135).

I took a look at the spectrum of the BS oplev error in pit and yaw with both loops enabled while the optic was stable. There is nothing alarmingly big except for some additional noise above 4Hz.

I have turned the BS oplev servo OFF for now.

 

  11199   Fri Apr 3 14:57:38 2015 manasaUpdateSUSBS oplev

The BS oplev has been misbehaving and kicking the optic from time to time since noon. The kicks are not strong enough to trip the watchdogs (current watchdog max counts for the sensors is 135).

I took a look at the spectrum of the BS oplev error in pit and yaw with both loops enabled while the optic was stable. There is nothing alarmingly big except for some additional noise above 4Hz.

I have turned the BS oplev servo OFF for now.

Attachment 1: BS_oplev_Apr3.png
BS_oplev_Apr3.png
  11198   Fri Apr 3 10:46:32 2015 SteveUpdateSUSOSEM sensor oscillation

ITMX, ETMY, BS and SRM are oscillating ?

Attachment 1: sensors1sec.png
sensors1sec.png
  11197   Fri Apr 3 05:17:36 2015 JenneUpdateLSCPRMI to 1f twelve times

I have 12 tick marks for times I got all the way to 1f for all 4 degrees of freedom in the PRFPMI.  The CARM / DARM transitions now succeed more than they fail, which is nice. 

At Q's suggestion, I am turning off all the violin filters in the MC2 path during the CARM transition.  This also means that I don't need any of the notches that Den and I put into the CARM_A and CARM_B filter banks last week, which were right at the edges of the violin notches.  Anyhow, this seems to make the transition much more likely to succeed.  I don't ever use the CM_SLOW FM10 "crossover helper" that Q had to use last night.  The violin filters are turned back on after the CARM transition is complete.  We don't ever need those other notches.

I checked the REFL165 vs. REFL55 transfer functions for PRCL and MICH, and they are mostly flat.  REFL55 seems like it'll give us extra phase for some reason.

I tried setting offsets for PRMI, but they seem to be strongly dependent on arm alignment.  I ended up being pretty confused, and since all the REFL signals are pretty close to zero (when CARM/DARM on RF, PRMI on 3f), I have given up on that avenue for tonight. 

I think many of my locklosses tonight (lost from the all 1f state) have been fast things, faster than the ADCs can handle.  On the lockloss plots that I've looked at, the FSS PC drive is railed at 10V about 200msec before I lose lock.  So, something (presumably in the fast CARM path) is making the MC/FSS loop unhappy.  I have plugged in the Agilent to the Out2 of the CM board, so that it looks at REFL11.  Unfortunately, this is after the input gain slider, so we don't see much until we're locked, but that seems fine.  A video camera is pointed at the screen, so that I get real time spectra.  It's hard to watch the TV at the same time as everything else, so I haven't witnessed the moment of lockloss in the fast spectrum yet.  Be careful when walking down the Yarm.  The tripod is partly in the walkway.

Q, I took a few TFs of the total CARM loop, although none of them are particularly good below a few kHz.  I can't push hard enough to get coherence, without blowing the lock.  TF data is in /users/jenne/PRFPMI/CM_TFs/CM_TFs_2Apr2015/ . 

I was worried for a while that, after I transition PRMI to 1f, I hear lots of low frequency rumbling.  However, watching the spectra (relative to references taken with CARM and DARM on RF, but PRMI on 3f), the low frequency error and control signals are staying the same for all 4 DoFs, but the high frequency for PRCL and MICH goes down significantly, so it's probably just that the low frequency stuff sounds more obvious, since it's not drowning in high frequency fuzz. 

 

  11196   Thu Apr 2 17:11:28 2015 ericqUpdateLSCNot much locking, Xover measurement

Whoops, I implemented the IOP downsampling filters wrong. Once I did that, it looked like just delay mismatch, so I added two more computation cycles for a total of four 16k cycles, which is maybe not so justified... Nevertheless, model and measurement now agree much better. Here are the corrected plots. 

 

Attachment 1: xOverModel.png
xOverModel.png
Attachment 2: loopModel.png
loopModel.png
  11195   Thu Apr 2 15:34:34 2015 ericqUpdateLSCNot much locking, Xover measurement

Here's the comparison of last night's crossover measurement to my loop model. Not stellar, but not totally off base. All of the digital filters are read directly from the foton filter file, and translated from their SOS coefficients, so they should be accurate. I may have tallied together the wrong arrangement of FMs, though. I will recheck. 

Although I don't have a measurement to compare it with yet (as I don't know where the crossover was, the filter statesolder, etc. for the older loop measurements), here's what my current CARM loop model looks like, just for kicks. Here, only the first CM board boost is on. If we turn on some super boosting, we can probably ease up on some of the digital boosts, lower the crossover frequency, and put some lowpass that suppress the violin filters' effect on the crossover and reduces digital sensing noise injection. 

Lastly, I'll just note that my current MIST model predicts that the CARM cavity pole should be at ~170Hz, and a peak arm transmission of 180 times single arm power. I saw powers of ~120 last night. 

Attachment 1: xOverModel.png
xOverModel.png
Attachment 2: loopModel.png
loopModel.png
  11194   Thu Apr 2 04:11:20 2015 ericqUpdateLSCNot much locking, Xover measurement

A paltry two locks tonight, but not entirely useless. I had some issues keeping the PRMI locked, which some additional boosts helped with. But, my feeling was that our crossover process is not tuned well. 

At full lock, both sub-loops have high gain around the crossover region, so the usual DTT loop transfer function measurement produces a meausrement of Gdigital/G_aopath (or minus that. I.e. I'm not currently 100% which is the bad phase in this plot, though it intuitive looks like 0 ). Thus, we can directly look at the crossover frequency and the effect of the different filters there. (I've also been working on an up-to-date CARM loop model today, so this will help inform that). 

Below, the black traces are the crossover at the end of the script when using the 120:500 "helper," and purple is without it. As we turn up the AO path gain, the trace "falls" from above, which explains why we can see instabilities around the violin filter. 

Having the helper on definitely made the probability of surviving the first overall CARM gain ramp higher, but it's not currently intuitively clear to me why that is the case. Afterwards, we can turn the helper off, to keep the shallower crossover shape. This is what I've put in to the up script for now. I also added a few seconds delay for when the script wants to switch DARM to RF only; I found it was maybe speeding too fast through this point.

DTT xml attached

Attachment 1: CARMxOver_Apr3.png
CARMxOver_Apr3.png
Attachment 2: Apr2_Xover.xml.zip
  11193   Thu Apr 2 01:45:44 2015 JenneUpdateLSCX Green Power drifting
Quote:

Have you tried a different set of laser temperatures?

Yep, that is how I got back to stable powers. 

  11192   Thu Apr 2 01:28:34 2015 JenneUpdateLSCX Green Power drifting

Have you tried a different set of laser temperatures?  I don't remember the value for the Xgreen, but whatever the value that matches PSL of 0.62ish and above seems to put the Xgreen laser at a bad temperature.  I think this is the mode-hopping region, and we sometimes lock to the wrong mode. 

So, FSS values of above 0.5ish are good, but they should be below 0.61ish. 

  11191   Wed Apr 1 23:56:36 2015 ericqUpdateLSCX Green Power drifting

Something funky is happening with the green light locked to the X arm. The green transmitted power is drifiting around. Maybe something weird is happening with the doubler? The digital thermal feedback loop is not on. 

The green has been locked on a TM00 mode this whole time. The step in power is me closing the PSL green shutter, but I'm not doing anything during the smooth changes in power. IR power is steady, so the alignment should be ok. I can't recover full power with the end PZT alignement either. 

 

Attachment 1: Xgreen_drifting.png
Xgreen_drifting.png
  11190   Wed Apr 1 16:52:02 2015 JenneUpdateLSCList of measurements

I'd like to get a concrete list of measurements written down, so that it's clear what needs to be done before I graduate.

Noise couplings:

  • Laser amplitude noise coupling to DARM
    • We don't have an AOM for ISS right now, but we should be able to just stick it back in the beam path, right?  I think Koji checked that the AOM was all okey-dokey recently.
    • AOM calibration should tell me how much the single-pass amplitude changes as a function of input signal.
  • Laser frequency noise coupling to DARM
    • Inject signal at the CM board input, which should be calibrated by looking at response to calibrated MC2 motion.
  • Marconi phase noise coupling to DARM
    • Marconi can produce internally or accept via BNC 0-10 rad of phase modulation.  Marconi spec sheet should give me rad/V for the input calibration.
  • Marconi amplitude noise coupling to DARM
    • Using external input of Marconi.  Marconi spec sheet should give me input calibration.
  • MICH err to DARM
    • Compare with Optickle model
  • PRCL err to DARM
    • Compare with Optickle model

Noise cancellation:

  • PRC angle
  • MICH noise removed from DARM (compare flat gain FF vs. Wiener FF)
  • PRCL noise removed from DARM (compare FF shape from model vs. Wiener FF)
  • MC length noise (equiv. to laser freq noise) removed from CARM & DARM

Are there things that I'm missing?  I've never had an IFO to characterize before.

  11189   Wed Apr 1 11:42:30 2015 manasaUpdateComputer Scripts / ProgramsPID script in python

Since none of us here are experts in pearl, I have put together a python script for a simple PID controller. This can be imported into any main scripts that will run the actual PID loop. The script, PID.py, exists in /scripts/general/

  11188   Wed Apr 1 10:50:59 2015 SteveUpdatePEMconstruction ahead

The 40m fenced area will start storing this large ~ 8000 lbs chamber on April 14. The asphalt will be cut, jack hammered the next 2-3 days in order to lay concrete.

Their schedule is from 8 to 5 starting tomorrow.  We are asking them to work from 6 to 3pm

ETMX is about 12-15 ft away

Attachment 1: largeMITchamber.jpg
largeMITchamber.jpg
  11187   Wed Apr 1 03:54:15 2015 JenneUpdateLSCPRMI to 1f twice

I got the PRMI transitioned from REFL165 over to REFL55 two times tonight.  Also, I had 2 long-ish locks, one 9 minutes, and one 6 minutes.  All the other locks were short - less than a minute or two.

I've done some shuffling around of the point in the CARM transition when the anti-boosts (1:20 filters) come on in the CARM A filter bank.  I've moved the turn-on of these filters a several gain steps earlier, but I'm not sure that they're in the best place yet.  Fiddling with the turn-on of the anti-boosts makes the big CARM oscillations last for longer or shorter - if they last too long, they blow the lock, so we don't want them to get too big.

The PRMI angular feedforward has helped a lot tonight, I think.  I've added a line to the up script to enable the output of the OAF after the PRMI is locked, and the down script turns it off again.  It's not so great when the PRM isn't aligned, since it's designed to work when the oplev is on, so it should be off unless the PRM is aligned.  I tried to get a comparison of off vs. on PRC powers with the arms resonating, but I can't hold the lock for long enough when the OAF is not on to get even one average on my 0.01Hz bandwidth spectrum. 

I've turned the arm ASC on a few times, but not every lock.  Around 12:34am, I set the offsets when CARM and DARM were on RF signals, and I had hand-aligned the ETMs to minimize the power at the AS port.  But, this wasn't a good spot for the next lock - the AS port was much darker with the ASC off for that lock.  It would be nice to think about trying some dither alignment, and then maybe resetting the setpoints every lock.  I'm using Q's original loop shapes, but as he left them yesterday, only actuating on the ETMs (with Yarm Yaw gain 0.7 rather than 0.9).

The CARM crossover might need more tuning.  There's some gain peaking around 400 Hz that goes mostly away if I turn the digital CARM gain down by 2dB.  (I'm not using any filters in the CM_SLOW filter bank).

I think that the CARM/DARM transition is more likely to be successful if the FSS slow DC is greater than 0.55ish.  So far this is pretty anecdotal, but I think I have more success when it's higher.  We should pay attention, and see if our trouble locking later in the nights correlates with smaller FSS slow DC values. 

I got the PRMI over to 1f two times, at 1:54am and at 2:25am.  I did not re-phase "POP"55 (which is the REFL55 signal), but I did check the values for the input matrix.  I needed MICH = 0.01*POP55Q and PRCL = 0.008*POP55I.  The first time I lost lock because I turned down the CARM digital gain too much.  The second time I forgot to turn down the PRCL gain (I was *actually* using 0.01*POP55I for the PRCL input matrix, but needed to lower the gain from -0.08 to -0.07, which is about the same as just using 0.008 in the input matrix).  Anyhow, I think PRCL loop oscillations were the cause of the second lockloss.


Here's a strip chart of my first lock of the night, which was the 9 minute lock.  Up until about -6 minutes, I was hand-aligning (including the dip around -7.5 minutes, where I was figuring out which direction to move the ETMs).  Around -3.5 minutes there is a significant dip down, that corrected itself.  By the time I realized that the power had gone down, and was trying to figure out why, it came back.  Maybe the same thing happened at the end of the lock, but it kept getting worse?  Self, re-look at this time (around 11:50pm) to find out why the power dips. 

My tummy feelings (without any data) make me think that this could be something with ITMX, like Q saw earlier today.  Or, maybe ETMX, like we've seen for ages.  Anyhow, my tummy feeling says this is an optic pointing problem.  I certainly think this might be the same thing we see at the end of many locks, the power going low suddenly.  So, it might give a big clue to our locklosses.  Maybe.

RXA: I've changed the above text into pink Comic Sans to lend it the appropriate level of gravitas, given its scientific justification.

Attachment 1: FirstLock_9min_arms150_1154pm.png
FirstLock_9min_arms150_1154pm.png
  11186   Tue Mar 31 22:27:43 2015 JenneUpdateModern ControlPreliminary PRMI angular Wiener results

Before locking for the evening, I wanted to try again implementing the Wiener filters that I had designed back in Jaunary (elog 10959). 

The problem then was that the newer version of Quack that I was using was doing weird things to me (elog 10993).  But, tonight I used the old quack3andahalf that we used to use for Wiener-related things, and that worked (for up to order 20 filters).  Actually, the pitch z-axis Wiener filter, when I copy the command string into Foton, says "Error" in the alternate box (the lower one).  I also get this error message if I try to put in filters that were greater than order 20, and have been split into several filters.  I'm not sure what's wrong, so for tonight I'm leaving out the pitch z-axis seismometer feed forward, and only using 20th order filters for all the rest.

So, pitch has feed forward signals from the T-240's x and y axes, and yaw has feed forward signals from all 3 seismometer channels.

At first, I just had the calculated Wiener filters, and a 10Hz lowpass, but the POP beam spot on the camera was getting slowly pushed away from the starting location.  So, I added a 0.01Hz cheby1 highpass filter, and that seems to have fixed that problem.  I need to go back to the simulations though, and see if this is going to cause extra noise to be injected (because of incorrect phase in the feed forward signal) at very low frequencies.  All 5 Wiener filter banks have a gain of -1.

I'm getting a factor of 4-5ish between 2Hz and 3Hz in both pitch and yaw.  What's interesting is that despite no direct angular suppression (as measured by the QPD) at higher frequencies, both POP22 and POPDC see improvement over a much broader range of frequencies.  I'll have to think about how to predict this RIN coupling in my budgets.

The time series data for these filters was collected 2 months ago, on the 29th of January.  So, it's nice to see that they work now too (although we have already seen that length feed forward signals are good over a many-month period).   

In uncalibrated units (I need to calibrate the QPD to microrad, and should probably quote the PD signals in RIN), here is the plot.  Blue trace (taken first) was with the feed forward on. Red trace (taken immediately afterward) was with feed forward off. This data is all PRMI-only, locked on REFL165 using Koji's recipe from elog 11174, including changing REFL165 phase to -14deg (from the -110 I found it at) for the no-arms case.

PRMI_31Mar2015.pdf

Attachment 1: PRMI_31Mar2015.pdf
PRMI_31Mar2015.pdf
  11185   Tue Mar 31 18:27:58 2015 ericqUpdateLSCSome locks
Quote:

Can we plot the arm power trend for multiple locks to see if it is associated with any thermal phenomenan in the IFO?

I'm currently more inclined to believe that the arm power trends have more to do with the arm alignments. Here's a 10 minute lock from last night, where the QPD servos were switched on about halfway through. I couldn't get Den's new servos to turn on without blowing the lock, so I reverted to my previous design, but still only actuated on the ETMs, with their oplevs still on. 

The most obvious feature is the reduction in power that seems to correspond to a ~10urad pitch deflection of ITMX when the lock begins. Is this optical spring action?

Also, it looks like the Y arm Yaw loop was badly tuned, and injecting noise. Ooops.


As of Den's QPD tuning, the QPD servos just actuate on the ETM. This next lock effectively had the QPD servos on the entire time, and we can see a similar drift in ITMX, and how ETMX then follows it to keep the QPD spot stationary. (Here, I'm plotting the QPD servo control signals, unlike above, so we can see X pitch servo output drift with the ITMX deflection)

Again, ITMX is moving in pitch by ~10urad when the interferometer starts resonating. If this is an optical spring, why does this just happen to ITMX? If it is digital shenanigans, how does it correlate with the lock, since there is nothing actuating on ITMX but oplevs and OSEM damping? Is light scattering into the ITMX OSEMs?

 

Attachment 1: qpdSwitch.png
qpdSwitch.png
Attachment 2: qpdAlways.png
qpdAlways.png
  11184   Tue Mar 31 09:03:32 2015 SteveUpdateVACRGA scan pd78 day 182

 

 

Attachment 1: RGApd78d182.png
RGApd78d182.png
  11183   Tue Mar 31 03:02:44 2015 KojiUpdateLSCSome locks

Assuming the carrier mode in PRC is stable and the SB is the one moving, can we just use the POP DC QPD to control PRM?

Can we plot the arm power trend for multiple locks to see if it is associated with any thermal phenomenan in the IFO?
They should be able to fit with an exp + DC.

  11182   Tue Mar 31 02:51:39 2015 ericqUpdateLSCSome locks

I had a handful of ~10 minute locks tonight. I intended to work on the 1f PRMI transition, but ended just familiarizing myself with the current scheme. 

Before touching anything, I committed the locking scripts to the svn. Unfortunately, the up script as I found it never worked for me tonight. I had to reintroduce the digitial crossover helper in CM_SLOW to get past the ramping up of the overall REFL11 gain. (With this is in place, there is some bad ringing around 200Hz for a time, but it goes away... or unlocks)

I did phase the PD formally known as REFL55 with an 800Hz PRM excitation while in full lock.42 to 102 degrees, ~30dB ratio between the I and Q peaks. However, come to think of it, how much does the CARM loop interfere with this?

The locklosses I had seemed to be due to a large fluctuation in all cavities' power. Maybe this will be helped by better PRC angular control, but we could maybe be helped by normalizing the digital part of the CARM loop by the arm transmissions once lock is acquired. 

  11181   Sat Mar 28 03:21:49 2015 denUpdateLSCtowards angular ff

Tonight I measured seismic noise coupling to beam spot on PR2. There is coherence of 0.9 from X to PIT and Y to YAW around the stack resonances. TF was fited using vectfit and put into static matrix of oaf in the elements T240X -> PRM PIT, T240Y -> PRM YAW. I think we should actuate on the error point of the PRM OL but I decided not to go for a model change tonight. Data from seismometers and POP QPD was obtained during the UTC time 04:06:00 - 04:50:00 when PRMI was locked on sideband

Interferometer was locking rather robustly and every lock lasted on the everage of 3 minutes. During these lock periods I incresed bandwidth of optical lever servos of BS and test masses from 4Hz up to 10Hz and then closed transmission QPD loops. It seems from the camera that lock losses correspond to strong motion of the beam on pop camera. Scripts that change OPLEV bandwidth are in /users/den "increase_ol_bandwidth.sh" "decrease_ol_bandwidth.sh". Script "engage_qpd_servos" turns off ETM oplevs and turns on ETM -> trans QPD servos. These scripts can be copied to locking directrly if are useful.

Please, note that transition from 3f to 1f should still be tuned. Only PRCL was stably controlled using 1f so far

  11180   Fri Mar 27 20:32:17 2015 KojiSummaryLSCLocking activity

- Adjutsed the IMC WFS operating point. The IMC refl is 0.42-0.43.

- The arms are aligned with ASS

- The X arm green was aligned with ASX. PZT offsets slides were adjusted to offload the servo outputs.

- I tried the locking once and the transition was successfull. I even tried the 3f-1f transition but the lock was lost. I wasn't sure what was the real cause.

I need to go now. I leave the IFO at the state that it is waiting for the arms locked with IR for the full locking trial.

  11179   Fri Mar 27 14:47:57 2015 KojiUpdateLSCals->pdh transition, prcl on 1f, alignment

Jenne and I interviewd Den this afternoon to make the things clear

- His "duty cycle" is not about the lengths of the lock stretch. He saids, the transition success probability is improved.

- For this improvement, the CARM transition procedure was modified to include turning on 1:20 (Z1P20) filter in CARM_A (i.e. ALS) once CARM_B (i.e. RF) dominates the loop in all frequency.

- I think this transition can be summarized like the attachment. At STEP4, the integration of the ALS is reduced. This actually does not change the stability of the servo as the servo stability is determined by the stability of the CARM_B loop. But this does further allow CARM_B to supress the noise. Or in other word, we can remove the noise coming from the CARM_A loop.

- The POP22 issue: Jenne has the trigger signal that is immune to this issue by adding some amount of POPDC for the trigger.
We can avoid the trigger issue by this technique. But if the issue is due to the true optical gain fluctuation, this may mean that the 11MHz optical gain is changing too much. This might be helped by PRC angular feedforward or RF 22MHz QPD at POP.

Attachment 1: CARM_transition.pdf
CARM_transition.pdf
  11178   Fri Mar 27 10:38:40 2015 steveUpdateVACvac rack UPS battery replaced

Batteries replaced after 3.5 years with Amstron AP-1250F2,  8x 12V 6Ah

Quote:

APC Smart -UPS 2200   model: SUA2200RM2U   batteries were replaced by compatible RBC43, 8x  12V5A

 

Attachment 1: UPSvac.jpg
UPSvac.jpg
  11177   Fri Mar 27 04:36:46 2015 denUpdateLSCals->pdh transition, prcl on 1f, alignment

Tonight I have modified transition steps from als to pdh signals. I have added 1:20 filters to CARM_A and DARM_A filter banks to make them unconditionally stable. These filters made locking more robust -- duty cycle is was ~70% tonight. I have also modified slow/ao crossover to avoid ringing up of lines above 1kHz.

Once AO is engaged with high bandwidth, REFL55 signal looks good and I transition PRCL from 165I to 55I. Optical gain compared to PRMI reduced from 55I/165I = -330 down to 55I/165I = 30 in full lock.

I worked on alignment of ETMs. Looking on the cameras I could improve arm power up to 160 and ifo visibility was 80%. POP22 fluctuated by ~50% and every few minutes we loose lock because POP22 almost touches zero.

  11176   Thu Mar 26 16:32:32 2015 JenneUpdateLSCREFL PDs get more light

Some more words on yesterday's REFL path work. 

The 90/10 BS that splits the light between REFL11 and REFL55 was placed back in August 2013, to compensate for the fact that REFL11 has a much larger RF transimpedance than REFL33.  See elog 9043 for details.

We had been operating for a long time with an embarrasingly small amount of light on the REFL PDs.  REFL11 used to have 80 uW, REFL33 used to have 400 uW and REFL55 used to have 700 uW.  REFL 165 was the only sane one, with about 15 mW of light.

After yesterday's work, the situation is now:

  Power incident [mW] PD responsivity [A/W] photocurrent [mA]
shot noise intercept
current [mA]
Ratio (photocurrent) /
(shot noise intercept current)
REFL 11 1.3 mW 0.7 0.91 mA 0.12 mA 7.6
REFL 33 13 mW 0.7 9.1 mA 0.52 mA 17.5
REFL 55 12 mW 0.7 8.4 mA 1.6 mA 5.3
REFL 165 50 mW 0.15 7.5 mA 1.06 mA 7.1

As an aside, I was foiled for a while by S vs. P polarizations of light.  The light transmitted through the PBS was P-pol, so the optics directing the beams to REFL11, 33 and 55 were all P-pol.  At first I completely removed the PBS and the waveplate, but didn't think through the fact that now my light would all be S-pol.  P-pol beam splitters don't work for S-pol (the reflection ratios are different, and it's just a terrible idea), so in the end I used the PBS to set the half waveplate so that all of my light was P-pol, and then removed the PBS but left the waveplate.  This means that all of the old optics are fine for the beams going to the 3 gold-box REFL PDs.  We don't have many S-pol beamsplitter options, so it was easier to use the waveplate to rotate the polarization. 

  11175   Thu Mar 26 10:41:06 2015 SteveUpdateIOOThe PMC is not clamped

The PMC is seated on 3 SS balls and it is free to move. I'm sure it will move in an earthquake. Not much, because the input and output K1 mirror frame will act as an earthquake stop.Atm2

Are there a touch of super glue on the balls? No, but there are V grooves at the bottom and on the top of each ball.Atm3

 

Attachment 1: IMG_0001.JPG
IMG_0001.JPG
Attachment 2: PMCstops.jpg
PMCstops.jpg
Attachment 3: PMCballv.jpg
PMCballv.jpg
  11174   Wed Mar 25 21:44:20 2015 KojiUpdateLSCIFO recovery / PRFPMI locking activity

[Koji, Den]

- Aligned the arms with ASS. It had alot of offset accumulated. We offloaded it to the suspension.

- We could lock the PRMIsb with the new setup.
PRCL: REFL165I (-14deg, analog +9dB)) -0.1, Locking FM4/5, Triggered FM 2
MICH: REFL165Q (-14deg, analog +9dB) -1.5, Locking FM4/5, Triggered FM2/6/9

- Demod phases at REFL were adjusted such that PRCL in Q signals were minimized :
REFL165 -80deg => -14deg
POP55 -63deg
REFL11 +164 => +7
REFL33 +136 => +133

Note: analog gains: REFL11: +18dB,  REFL33: +30dB, POP55: +12dB, REFL165: +9dB

- Try some transition between REFL signals to check the signal quality.
Measure TFs between the REFL signals

PRCL gain
REFL11I/REFL165I = +58
REFL33I/REFL165I = +8.5
POP55I /REFL165I = -246

MICH gain
REFL11Q/REFL165Q = +11
REFL33Q/REFL165Q = -1.5
POP55Q /REFL165Q = +280

- This resulted us to figure out the relationships of the numbers in the input matrix 

REFL55I/Q -4e-3/4e-3
REFL165I/Q 1.0/1.0 (reference)
REFL11I/Q  0.02/0.1
REFL33I/Q +0.12/-0.7


Full locking trial

Arm locked -> ALS -> Arm offset locked
PRMI locking
REFL165 phase tuned -110deg
PRCL gain -0.1 / MICH gain -2

We needed script editing.
Previous script saved in: /opt/rtcds/caltech/c1/scripts/PRFPMI/carm_cm_up_BACKUP.sh

Change:
- PRMI gain setting (input matrix & servo gain)
- CARM/DARM transition setting (see below)


The current CARM/DARM transition procedure:

== CARM TRANSITION (PART1) ==
- CM REFL1 gain is set to be -32
- CARM_B is engaged and the gain is ramped from 0 to +2.5
- Turn on FM7 (integrator)
- MC IN2 (AO path) engaged
- MC IN2 gain increased from -32 to -21

== DARM TRANSITION (PART1) ==
- DARM_B is engaged and the gain is ramped from 0 to +0.1
- Turn on FM7 (integrator)

== CARM TRANSITION (PART2) ==
- CM REFL1 Gain is increased from -32 to -18
- Ramp down CARM A gain to 0

== DARM TRANSITION (PART2) ==
- DARM_B gain is incrased to 0.37. At the same time DARM_A gain is reduced to 0


We succeeded to make the transition several times in the new setting.

- But later the transition got hard. We started to see big jump of the arm trans (TRX/Y 50->100) at the CARM transition.

- We tested the PRCL transition from 165MHz to 55MHz. 55MHz (i.e. POP55 which is REFL55PD) looks alot better now.

- ~1:30 The PMC was realigned. This  increased PMC_TRANS about 10%. This let the Y arm trans recover ~1.00 for the single arm locking

- Decided to end around 3:00AM

  11173   Wed Mar 25 18:48:11 2015 KojiSummaryLSC55MHz demodulators inspection

[Koji Den EricG]

We inspected the {REFL, AS, POP}55 demodulators.

Short in short, we did the following changes:

- The REFL55 PD RF signal is connected to the POP55 demodulator now.
Thus, the POP55 signals should be used at the input matrix of the LSC screens for PRMI tests.

- The POP55 PD RF signal is connected to the REFL55 demodulator now.

- We jiggled the whitening gains and the whitening triggers. Whitening gains for the AS, REFL, POP PDs are set to be 9, 21, 30dB as before.
However, the signal gain may be changed. The optimal gains should be checked through the locking with the interferometer.


- Test 1

Inject 55.3MHz signal to the demodulators. Check the amplitude in the demodulated signal with DTT.
The peak height in the spectrum was calibrated to counts (i.e. it is not counts/rtHz)
We check the amplitude at the input of the input filters (e.g. C1:LSC-REFL55_I_IN1). The whitening gains are set to 0dB.
And the whitening filters were turned off.

REFL55
f_inj = 55.32961MHz -10dBm
REFL55I @999Hz  22.14 [cnt]
REFL55Q @999Hz  26.21 [cnt]


f_inj = 55.33051MHz -10dBm
REFL55I @ 99Hz  20.26 [cnt]  ~200mVpk at the analog I monitor
REFL55Q @ 99Hz  24.03 [cnt]


f_inj = 55.33060MHz -10dBm
REFL55I @8.5Hz  22.14 [cnt]
REFL55Q @8.5Hz  26.21 [cnt]


----
f_inj = 55.33051MHz -10dBm
AS55I   @ 99Hz 585.4 [cnt]
AS55Q   @ 99Hz 590.5 [cnt]   ~600mVpk at the analog Q monitor

f_inj = 55.33051MHz -10dBm
POP55I  @ 99Hz 613.9 [cnt]   ~600mVpk at the analog I monitor
POP55Q  @ 99Hz 602.2 [cnt]

We wondered why the REFL55 has such a small response. The other demodulators seems to have some daughter board. (Sigg amp?)
This maybe causing this difference.

-----

- Test 2

We injected 1kHz 1Vpk AF signal into whitening board. The peak height at 1kHz was measured.
The whitening filters/gains were set to be the same condition above.

f_inj = 1kHz 1Vpk
REFL55I 2403 cnt
REFL55Q
2374 cnt
AS55I   2374 cnt
AS55Q   2396 cnt
POP55I  2365 cnt
POP55Q
  2350 cnt

So, they look identical. => The difference between REFL55 and others are in the demodulator.

  11172   Wed Mar 25 18:46:14 2015 JenneUpdateLSCREFL PDs get more light

After discussions during the meeting today, I removed the PBS from the REFL path, which gives much more light to REFL11, REFL33 and REFL55.  Also, the ND1.5 in front of REFL165 was replaced with ND1.1, so that REFL165 now gets 50mW of light.  REFL11 gets about 1.3mW, REFL33 gets about 13mW and REFL55 gets about 12mW. 

No locking, and importantly no re-phasing of any PDs has been done yet. 

Here is an updated diagram of the REFL branching ratios.

Attachment 1: AS_REFL_branchingRatios_25Mar2015.png
AS_REFL_branchingRatios_25Mar2015.png
  11171   Wed Mar 25 18:27:34 2015 KojiSummaryGeneralSome maintainance

- I found that the cable for the AS55 LO signal had the shielding 90% broken. It was fixed.

- The Mon5 monitor in the control room was not functional for months. I found a small CRT down the east arm.
It is now set as MON5 showing the picture from cameras. Steve, do we need any safety measure for this CRT?

  11170   Wed Mar 25 08:29:31 2015 SteveUpdateLSCMeh

  I turned on the HEPA at the south end during the LSC. Sorry I ment to turn it off.

Quote:

[Jenne, Den]

Overall a "meh" night for locking I think.  The script to all-RF worked several times earlier in the evening, although it was delicate and failed at least 50% of the time.  Later in the evening, we couldn't get even ~10% of the lock attempts all the way to RF-only. 

Den looked into angular things tonight.  With the HEPA bench at the Xend on (which it was found to be), the ETMX oplevs were injecting almost a factor of 10 noise (around 10ish Hz?) into the cavity axis motion (as seen by the trans QPD) as compared to oplevs off.  Turning off the HEPA removed this noise injection.

Den retuned the QPD trans loops so that they only push on the ETMs, so that we can turn off the ETM oplevs, and leave the ITMs and their oplevs alone. 

We are worried again about REFL55.  There is much more light on REFL55 than there is on REFL11 (a 90/10 beam splitter divides the light between them), and we see this in the DC output of the PDs, but there seems to be very little actual signal in REFL55.  Den drove a line (in PRCL?) while we had the PRMI locked with the arms held off resonance, and REFL55 saw the line a factor of 1,000 less than REFL 11 or REFL165.  The analog whitening gain for REFL11 is +18dB, and for REFL55 is +21dB, so it's not that we have significantly less analog gain (that we think).  We need to look into this tomorrow.  As of now, we don't think there's much hope for transitioning PRMI to REFL55 without a health checkup. 

 

  11169   Wed Mar 25 03:31:18 2015 JenneUpdateLSCMeh

[Jenne, Den]

Overall a "meh" night for locking I think.  The script to all-RF worked several times earlier in the evening, although it was delicate and failed at least 50% of the time.  Later in the evening, we couldn't get even ~10% of the lock attempts all the way to RF-only. 

Den looked into angular things tonight.  With the HEPA bench at the Xend on (which it was found to be), the ETMX oplevs were injecting almost a factor of 10 noise (around 10ish Hz?) into the cavity axis motion (as seen by the trans QPD) as compared to oplevs off.  Turning off the HEPA removed this noise injection.

Den retuned the QPD trans loops so that they only push on the ETMs, so that we can turn off the ETM oplevs, and leave the ITMs and their oplevs alone. 

We are worried again about REFL55.  There is much more light on REFL55 than there is on REFL11 (a 90/10 beam splitter divides the light between them), and we see this in the DC output of the PDs, but there seems to be very little actual signal in REFL55.  Den drove a line (in PRCL?) while we had the PRMI locked with the arms held off resonance, and REFL55 saw the line a factor of 1,000 less than REFL 11 or REFL165.  The analog whitening gain for REFL11 is +18dB, and for REFL55 is +21dB, so it's not that we have significantly less analog gain (that we think).  We need to look into this tomorrow.  As of now, we don't think there's much hope for transitioning PRMI to REFL55 without a health checkup. 

  11168   Tue Mar 24 18:47:10 2015 ericqUpdateLSCAO Path engaged

Jenne has more detailed notes about how things went down last night, but I figure I should write about how we got the AO path stably up. 

As the carm_cm_up script stood after Jenne and Den's work last week, the CARM loop looked like the gold trace in the loop shape plot I posted in the previous elog. The phase bubble was clearly enlarged by the AO path, but there was some bad crossover instability brewing at 400 Hz. This was evident as a large noise peak, and would lead to lock loss if we tried to increase the overall CARM gain.

Quote:

 

As with our single arm CM board locking adventures, it was useful to have a filter that made the digital loop shape steeper around the crossover region, so that the 1/f AO+cavity pole shape played nice with the digital slope. As in the single arm trials, this effectively meant undoing the cavity pole compensating zero with a corresponding pole, letting the physical cavity pole do the steepening. This is only possible once the AO path has bestowed some phase upon you. A zero at a somewhat higher frequency (500Hz) gives the digital loop back some phase, which is neccesary to stay locked when the loop has only a few hundred Hz UGF, and the digital phase still matters. This gives us the purple trace. 

This provided us with a loop shape that could smoothly be ramped up in overall gain towards UGFs of multiple kHz (red trace). At this point we could reliably turn on the first boost, which will help in transitioning the PRMI to 1f signals (green trace). We didn't want to ramp it up too much, as we saw that the phase bubble likely ended not much higher than 100kHz, and the OLG magnitude was flattening pretty clearly around 40kHz. While we could turn on a super boost, it didn't look too nice, as we would have to stay at low phase margin to avoid bad gain peaking (blue trace).

As could be seen in the noise spectra that Jenne showed, you can see the violin notches in the CARM noise. This means we are injecting the digital loop noise all over the place. We attempted rolling off the digital loop (by undoing the zero at 500Hz), but found this made the gain at ~200Hz crash down, almost becoming unstable. We likely haven't positioned the crossover frequency in the ideal place for doing this. 

We didn't really give the interferometer any time to see how the long term stability was, since we wanted to poke around and measure as much as we could. While not every attempt would get us all the way there, the current carm_cm_up's success rate at achieving multi-kHz CARM bandwidth was pretty good (probably more than 50%) and the whole thing is still pretty snappy. 

ELOG V3.1.3-